# Optimistic Estimation of Convergence in Markov Chains with the Average-Mixing Time

### v2

@@ 5   Proofs
$\begin{split}\mathcal{B}_{p}^{(s)}&\leq\beta_{0}^{1/p}+\beta(s)^{1/p}\sum_{t=1}^{\lfloor(n-1)/s\rfloor-2}\frac{1}{t^{b/p}}\leq\beta_{0}^{1/p}+\zeta(b/p)\beta(s)^{1/p}.\\ \end{split}$
5.5   Proof of Theorem 3.2
Recall that for a uniformly ergodic Markov chain,
$\beta(s)\leq d(s)\leq d(\lfloor s/t_{\mathsf{mix}}\rfloor t_{\mathsf{mix}})\leq\exp(-\lfloor s/t_{\mathsf{mix}}\rfloor\log 2)\leq 2\exp\left(-\frac{\log 2}{t_{\mathsf{mix}}}s\right),$
where the third inequality is adapted from Levin et al. (2009, (4.33)), i.e. $\beta$ is decays exponentially with $\beta_{0}=2$ and $\beta_{1}=\log(2)/t_{\mathsf{mix}}$. Thus, from the fact that $x>0\implies 1/(1-\exp(-x))\leq 1+1/x,$ and Corollary 3.1, it holds that
$\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq 4\sqrt{\frac{2^{1/p}}{\lfloor(n-1)/s\rfloor}\left(1+\frac{pt_{\mathsf{mix}}}{s\log 2}\right)\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}},\\ \end{split}$
and $(i)$ follows from $n>s+1$ and $x>1\implies\lfloor x\rfloor\geq x/2$ and $2^{1/p}\leq 2$. From $(i)$, when
$n\geq 1+\frac{2^{8}}{\varepsilon^{2}}\left(s+\frac{p}{\log 2}t_{\mathsf{mix}}\right)\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p},$
@@
$\mathbb{P}_{\pi}\left(\left|\widehat{\beta}(s)-\beta(s)\right|>\varepsilon\right)\leq\mathbb{P}_{\pi}\left(\left|\left|\widehat{\beta}(s)-\beta(s)\right|-\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|\right|>\varepsilon/2\right).$
@@
$x_{[s]}^{n}=x_{1},x_{1+s},x_{1+2s},\dots,x_{1+\lfloor(n-1)/s\rfloor s}\in\mathcal{X}$
be some $s$-skipped trajectory, let $t\in\left\{0,\dots,\lfloor(n-1)/s\rfloor\right\}$, $x^{\prime}_{1+ts}\in\mathcal{X}$, and denote
$x^{n,(t)}=x_{1},x_{1+s},x_{1+2s},\dots,x_{1+(t-1))s},x^{\prime}_{1+ts},x_{1+(t+1)s}\dots,x_{1+\lfloor(n-1)/s\rfloor s}$
the trajectory where the observation $x_{1+ts}$ at time $1+ts$ was replaced with observation $x^{\prime}_{1+ts}$. We show that the estimator satisfies the bounded differences condition. We decompose,
$\begin{split}&\left|\left|\widehat{\beta}(s)(X)-\beta(s)\right|-\left|\widehat{\beta}(s)(X^{(t)})-\beta(s)\right|\right|\\ &\leq\left|\widehat{\beta}(s)(X)-\widehat{\beta}(s)(X^{(t)})\right|\\ &\leq\frac{1}{\lfloor(n-1)/s\rfloor}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{xx^{\prime}}^{(s)}(X)-N_{xx^{\prime}}^{(s)}(X^{(t)})\right|\\ &+\frac{1}{\lfloor(n-1)/s\rfloor^{2}}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{x}^{(s)}(X)N_{x^{\prime}}^{(s)}(X)-N_{x}^{(s)}(X^{(t)})N_{x^{\prime}}^{(s)}(X^{(t)})\right|.\\ \end{split}$
@@
$\begin{split}&\frac{1}{\lfloor(n-1)/s\rfloor^{2}}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{x}^{(s)}(X)N_{x^{\prime}}^{(s)}(X)-N_{x}^{(s)}(X^{(t)})N_{x^{\prime}}^{(s)}(X^{(t)})\right|\\ &\leq\frac{1}{\lfloor(n-1)/s\rfloor^{2}}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{x}^{(s)}(X)N_{x^{\prime}}^{(s)}(X)-N_{x}^{(s)}(X)N_{x^{\prime}}^{(s)}(X^{(t)})\right|\\ &+\frac{1}{\lfloor(n-1)/s\rfloor^{2}}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{x}^{(s)}(X)N_{x^{\prime}}^{(s)}(X^{(t)})-N_{x}^{(s)}(X^{(t)})N_{x^{\prime}}^{(s)}(X^{(t)})\right|\\ &\leq\frac{1}{\lfloor(n-1)/s\rfloor}\sum_{x^{\prime}\in\mathcal{X}}\left|N_{x^{\prime}}^{(s)}(X)-N_{x^{\prime}}^{(s)}(X^{(t)})\right|+\frac{1}{\lfloor(n-1)/s\rfloor}\sum_{x\in\mathcal{X}}\left|N_{x}^{(s)}(X)-N_{x}^{(s)}(X^{(t)})\right|.\\ \end{split}$
@@
$\left|\left|\widehat{\beta}(s)(X)-\beta(s)\right|-\left|\widehat{\beta}(s)(X^{(t)})-\beta(s)\right|\right|\leq\frac{4}{\lfloor(n-1)/s\rfloor}+\frac{2}{\lfloor(n-1)/s\rfloor}+\frac{2}{\lfloor(n-1)/s\rfloor}.$
@@
which completes the proof of claim $(ii)$. ∎
5.6   Proof of Theorem 3.3
Suppose that
$n-1\geq\frac{2^{9}}{\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon)\max\left\{\mathcal{J}_{p},18\log\left(\frac{9\cdot 2^{10}}{\delta\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon)\right)\right\}.$
Invoking Lemma 5.1 with
$a\doteq\frac{9\cdot 2^{10}}{\delta\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon),\qquad b\doteq\frac{2}{\delta_{n-1}},$
and where $\delta_{n-1}\doteq\frac{\delta}{n-1}$, it holds that
$n-1\geq\frac{2^{9}}{\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon)\max\left\{\mathcal{J}_{p},9\log\frac{2}{\delta_{n-1}}\right\}.$ (14)
Writing $S=S(n,\varepsilon,\delta,p,\mathcal{J}_{p})$, we thus have
$\begin{split}n-1&\geq\frac{n-1}{2}+\frac{2^{8}}{\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon)\max\left\{\mathcal{J}_{p},9\log\frac{2}{\delta_{n-1}}\right\}\\ &\geq\frac{2^{8}}{\varepsilon^{2}}\left[S+\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon)\right]\max\left\{\mathcal{J}_{p},9\log\frac{2}{\delta_{n-1}}\right\}\\ &\geq\frac{2^{8}}{\varepsilon^{2}}\max\left\{\left(S+\frac{p}{\log 2}t_{\mathsf{mix}}\right)\mathcal{J}_{p},9(S+t_{\mathsf{mix}})\log\frac{2}{\delta_{n-1}}\right\}.\end{split}$
Furthermore, observing from (14) that
$S\geq\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon)\geq t^{\sharp}_{\mathsf{mix}}(\varepsilon),$
it follows that $\beta(s)<\varepsilon$ for $s>S$. An application of the union bound and Theorem 3.2$(ii)$ yield that
$\mathbb{P}_{\pi}\left(\left\|\widehat{\beta}-\beta\right\|_{\infty}>\varepsilon\right)\leq\sum_{s=1}^{S}\mathbb{P}_{\pi}\left(\left|\widehat{\beta}(s)-\beta(s)\right|>\varepsilon\right)\leq S\delta_{n-1}\leq\delta.$
The claim then follows from $t_{\mathsf{mix}}(\varepsilon)\leq\lceil\log_{2}\varepsilon^{-1}\rceil t_{\mathsf{mix}}$ for any $\varepsilon\in(0,1/4)$ (Levin et al., 2009, (4.34)). ∎
5.7   Proof of Theorem 3.4
We first bound the probability that the estimator overshoots, that is $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)>t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))$. On this event, for any $s\leq t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))$, and in particular for $s=t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))$, it must be that $\widehat{\beta}(s)>\xi$. From Theorem 3.2$(ii)$, for
$n\geq 1+\frac{2^{8}}{\xi^{2}\varepsilon^{2}}\max\left\{\left(s+\frac{p}{\log 2}t_{\mathsf{mix}}\right)\left\|Q^{(s)}\right\|^{1-1/p}_{1-1/p},9(s+t_{\mathsf{mix}})\log\frac{4}{\delta}\right\},$
@@
As a second step, we bound the probability that $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)$ underestimates the true parameter, that is $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)<t_{\mathsf{mix}}^{\sharp}(\xi(1+\varepsilon))$. From Theorem 3.2$(ii)$ and an application of the union bound, when for any $s\leq t^{\sharp}_{\mathsf{mix}}(\xi(1+\varepsilon))$ the trajectory length satisfies
$n\geq 1+\frac{2^{8}}{\xi^{2}\varepsilon^{2}}\max\left\{\left(s+\frac{p}{\log 2}t_{\mathsf{mix}}\right)\left\|Q^{(s)}\right\|^{1-1/p}_{1-1/p},9(s+t_{\mathsf{mix}})\log\frac{4t^{\sharp}_{\mathsf{mix}}(\xi(1+\varepsilon))}{\delta}\right\},$

### v4

@@ Appendix C Proofs
$\begin{split}\mathcal{B}_{p}^{(s)}&\leq\beta_{0}^{1/p}+\beta_{0}^{1/p}\sum_{t=1}^{\lfloor(n-1)/s\rfloor-1}\exp\left(-\beta_{1}(st)^{b}/p\right)\\ &\leq\beta_{0}^{1/p}+\beta_{0}^{1/p}\exp(-\beta_{1}s^{b}/p)+\beta_{0}^{1/p}\sum_{t=2}^{\infty}\exp\left(-\beta_{1}s^{b}t^{b}/p\right)\\ &\leq\beta_{0}^{1/p}+\beta_{0}^{1/p}\exp(-\beta_{1}s^{b}/p)+\beta_{0}^{1/p}\int_{1}^{\infty}\exp(-\beta_{1}s^{b}t^{b}/p)dt\\ &=\beta_{0}^{1/p}+\beta_{0}^{1/p}\exp(-\beta_{1}s^{b}/p)+\beta_{0}^{1/p}\left[\frac{-1}{b(\beta_{1}s^{b}/p)^{1/b}}\Gamma\left(\frac{1}{b},\beta_{1}s^{b}t^{b}/p\right)\right]_{1}^{\infty}.\end{split}$
@@
C.5  Proof of Theorem 3.2
We first focus on the estimation of $\beta(1)$. The general result for $\beta(s)$ will be obtained by considering an $s$-skipped chain. First, similar to the proof of Theorem 3.1, we decompose using the triangle inequality,
$\begin{split}2\left|\widehat{\beta}(1)-\beta(1)\right|&\leq\sum_{x,x^{\prime}\in\mathcal{X}}\left|\frac{N_{xx^{\prime}}}{n-1}-Q(x,x^{\prime})\right|+2\sum_{x\in\mathcal{X}}\left|\frac{N_{x}}{n-1}-\pi(x)\right|.\end{split}$
We treat the two sums separately,
$\mathbb{P}_{\pi}\left(\left|\widehat{\beta}(1)-\beta(1)\right|>\varepsilon\right)\leq\mathbb{P}_{\pi}\left(\sum_{x,x^{\prime}\in\mathcal{X}}\left|\frac{N_{xx^{\prime}}}{n-1}-Q(x,x^{\prime})\right|>\varepsilon\right)+\mathbb{P}_{\pi}\left(\sum_{x\in\mathcal{X}}\left|\frac{N_{x}}{n-1}-\pi(x)\right|>\varepsilon/2\right).$ (17)
Let us focus on the second probability statement. For simplicity, we assume that n1=2Brn-1=2Br, where we decompose the process into $2B$ Bernstein blocks (refer to Figure 4) of size $r$.
[…]
Therefore, applying Yu (1994, Corollary 2.7),
$\mathbb{P}_{\pi}\left(\sum_{x\in\mathcal{X}}\left|\frac{N_{x}}{n-1}-\pi(x)\right|>\varepsilon/2\right)\leq\sum_{\sigma\in\{0,1\}}\mathbb{P}_{\pi}\left(\frac{1}{B}\sum_{b=1}^{B}Z_{b}^{(\sigma)}>\varepsilon/2\right)+(B-1)\beta(r),$ (18)
where
$Z_{b}^{(\sigma)}\doteq\frac{1}{r}\sum_{x\in\mathcal{X}}\left|\sum_{t=1}^{r}\left(\delta\left[X_{t}^{[2b-\sigma]}=x\right]-\pi(x)\right)\right|$
are independent random variables. First, observing that $Z_{b}^{(\sigma)}$ is a function of a Markov chain $X_{1},\dots,X_{r}$ stationarily sampled from $P$, we upper bound the expected value as in Theorem 3.1,
$\begin{split}\mathbb{E}_{\pi}Z_{b}^{(\sigma)}&\stackrel{{\scriptstyle(i)}}{{\leq}}\frac{1}{r}\sum_{x\in\mathcal{X}}\sqrt{\operatorname{Var}_{\pi}\left[\sum_{t=1}^{r}\delta\left[X_{t}^{[2b-\sigma]}=x\right]\right]}\\ &\stackrel{{\scriptstyle(ii)}}{{\leq}}\frac{2}{\sqrt{r}}\left\|\pi\right\|_{(1-1/p)/2}^{(1-1/p)/2}\sqrt{\mathcal{B}_{p}},\end{split}$
where $(i)$ is Jensen’s inequality and $(ii)$ is an application of Lemma C.2. Thus for
$r\geq\frac{64\left\|\pi\right\|_{(1-1/p)/2}^{1-1/p}\mathcal{B}_{p}}{\varepsilon^{2}},$
it holds that
$\mathbb{E}_{\pi}Z_{b}^{(\sigma)}\leq\varepsilon/4,$
and it remains to bound
$\mathbb{P}_{\pi}\left(\frac{1}{B}\sum_{b=1}^{B}Z_{b}^{(\sigma)}-\mathbb{E}_{\pi}Z_{b}^{(\sigma)}>\varepsilon/4\right).$
The random variables $Z_{b}^{(\sigma)}$ are bounded. Indeed,
[…]
Therefore, from an application of Hoeffding’s inequality,
$\mathbb{P}_{\pi}\left(\frac{1}{B}\sum_{b=1}^{B}Z_{b}^{(\sigma)}-\mathbb{E}_{\pi}Z_{b}^{(\sigma)}>\varepsilon/4\right)\leq\exp\left(-\frac{2(B\varepsilon/4)^{2}}{B2^{2}}\right)=\exp\left(-\frac{B\varepsilon^{2}}{32}\right).$
Plugging in (18),
$\mathbb{P}_{\pi}\left(\sum_{x\in\mathcal{X}}\left|\frac{N_{x}}{n-1}-\pi(x)\right|>\varepsilon/2\right)\leq 2\exp\left(-\frac{B\varepsilon^{2}}{32}\right)+(B-1)\beta(r).$
For $B\geq\frac{32\log(8/\delta)}{\varepsilon^{2}}$ and $r\geq t^{\sharp}_{\mathsf{mix}}(\xi(\varepsilon,\delta))$ with
$\xi(\varepsilon,\delta)=\frac{\varepsilon^{2}\delta}{128\log(8/\delta)},$
the above probability is smaller than $\delta/2$ In other words, regardless of uniform ergodicity, for
$n\geq 1+\frac{64\log(8/\delta)}{\varepsilon^{2}}\max\left\{\frac{64}{\varepsilon^{2}}\left\|\pi\right\|_{(1-1/p)/2}^{1-1/p}\mathcal{B}_{p},t^{\sharp}_{\mathsf{mix}}(\xi(\varepsilon,\delta))\right\},\qquad\xi(\varepsilon,\delta)=\frac{\varepsilon^{2}\delta}{128\log(8/\delta)},$
it holds that
$\mathbb{P}_{\pi}\left(\sum_{x\in\mathcal{X}}\left|\frac{N_{x}}{n-1}-\pi(x)\right|>\varepsilon/2\right)\leq\delta/2.$
A similar approach for the sliding-window chain $(X_{t},X_{t+1})_{t\in\mathbb{N}}$ leads to a bound on the first term of (17). Finally, considering the $s$-skipped version of a chain, and noting that for any $\xi\in(0,1)$, ${t^{\sharp}_{\mathsf{mix}}}^{(s)}(\xi)\leq\lceil t^{\sharp}_{\mathsf{mix}}(\xi)/s\rceil$ yields the claim for $\beta(s)$. ∎
C.6  Proof of Theorem 3.3
Our starting point for this proof if (16). However, instead of Rio’s inequality, we rely on techniques in Paulin (2015) to obtain an upper bound on the two variance terms. The mixing time of the $s$-skipped pairs of observations is upper bounded by $t_{\mathsf{mix}}^{(s)}+1\leq\lceil t_{\mathsf{mix}}/s\rceil+1$ (Wolfer and Kontorovich, 2021, Lemma 6.1). Combining Paulin (2015, Theorem 3.2, Proposition  3.4), it follows that
[…]
which yields
[…]
The claim with $C_{\mathsf{u.erg.}}\leq 3\sqrt{2}$ follows from $n>s+1$ and $x>1\implies\lfloor x\rfloor\geq x/2$ and $s(\lceil t_{\mathsf{mix}}/s\rceil+1)\leq t_{\mathsf{mix}}+2s$.
C.7  Proof of Theorem 3.4
From Theorem 3.3, when
$n\geq 1+\frac{4C_{\mathsf{u.erg.}}^{2}}{\varepsilon^{2}}(t_{\mathsf{mix}}+2s)\mathcal{J}_{\infty}^{(s)},$
@@
[…]
@@
$\bar{x}=(x_{1},x_{2},\dots,x_{n})\in\mathcal{X}^{n},$
and for $t\in[n]$ and $x^{\prime}_{t}\in\mathcal{X}$, let us denote
$\bar{x}^{\left\{t\right\}}\doteq(x_{1},x_{2},\dots,x_{t-1},x^{\prime}_{t},x_{t+1}\dots,x_{n})$
the trajectory where the observation $x_{t}$ at time $t$ was replaced with observation $x^{\prime}_{t}$. We show that the absolute deviation $\left|\widehat{\beta}(s)-\beta(s)\right|\colon\mathcal{X}^{n}\to\mathbb{R}$ satisfies the bounded differences condition. We decompose,
[…]
@@
[…]
@@
$\left|\left|\widehat{\beta}(s)(\bar{x})-\beta(s)\right|-\left|\widehat{\beta}(s)(\bar{x}^{\left\{t\right\}})-\beta(s)\right|\right|\leq\frac{4}{\lfloor(n-1)/s\rfloor}+\frac{2}{\lfloor(n-1)/s\rfloor}+\frac{2}{\lfloor(n-1)/s\rfloor}.$
@@
which completes the proof. ∎
C.8  Proof of Lemma 3.2
We first bound the probability that the estimator overshoots the target interval, that is $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)>t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))$. On this event, for any $s\leq t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))$, and in particular for $s=t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))$, it must be that $\widehat{\beta}(s)>\xi$. From our assumption
$n\geq n_{0}(P,t^{\sharp}_{\mathsf{mix}}(\xi(1-\varepsilon)),\xi\varepsilon,\delta/2),$
@@
As a second step, we bound the probability that $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)$ underestimates the true parameter, that is $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)<t_{\mathsf{mix}}^{\sharp}(\xi(1+\varepsilon))$. From our assumption and an application of the union bound, when for any $s\leq t^{\sharp}_{\mathsf{mix}}(\xi(1+\varepsilon))$ the trajectory length satisfies
$n\geq n_{0}\left(P,s,\xi\varepsilon,\frac{\delta}{2t^{\sharp}_{\mathsf{mix}}(\xi)}\right),$